OpenStudy (anonymous):

Identify the 17th term of a geometric sequence where a1 = 16 and a5 = 150.06. Round the common ratio and 17th term to the nearest hundredth.

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

so in this one a is 16 150.06=16*r^4

OpenStudy (anonymous):

r = 1.749992711

OpenStudy (anonymous):

what does the r stand for?

OpenStudy (anonymous):

so the 17 term is 16 * (1.749992711^16) = 123 794.064

OpenStudy (anonymous):

r stands for ratio. but it is arbitrary tho.

OpenStudy (anonymous):

a17 ≈ 123,802.31 a17 ≈ 30,707.05 a17 ≈ 19,684.01 a17 ≈ 216,654.05

OpenStudy (anonymous):

actually if your round r r = 1.74999271 r = 1.75 so 16 * (1.75^16) = 123 802.31

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

That was correct! Ty!!!